# Competing skin effect and quasiperiodic localization in the non-Hermitian Su–Schrieffer–Heeger chain: Reentrant delocalization, spectral topology destruction, and entanglement suppression

### 20 Mar 2026 v1

@@ I Introduction
The Su–Schrieffer–Heeger (SSH) model [1] is the prototypical one-dimensional topological insulator, supporting symmetry-protected zero-energy edge states when the intercell hopping exceeds the intracell hopping. Adding an Aubry–André–Harper (AAH) on-site potential introduces a metal–insulator transition at a critical modulation strength $\lambda_{c}=2w$, above which all single-particle states become localized [2]. Separately, introducing nonreciprocal hopping renders the Hamiltonian non-Hermitian and gives rise to the non-Hermitian skin effect (NHSE)—the anomalous localization of an extensive number of eigenstates at a single boundary under open boundary conditions [4, 5, 6]. The NHSE is intimately tied to point-gap topology in the complex energy spectrum [7, 8]: the number of times the periodic-boundary-condition spectrum winds around a base energy determines the existence and direction of skin accumulation.
Each pairwise combination has been studied: the Hermitian SSH–AAH model [17, 18], the non-Hermitian AAH model [11, 12, 13], and the non-Hermitian SSH model without disorder [4, 5]. The non-Hermitian AAH model, in particular, exhibits a competition between skin-effect localization and quasiperiodic localization, with a localization–skin-effect phase transition studied both theoretically [11, 12, 14] and experimentally via photonic quantum walks [16]. However, the combined model—a dimerized (SSH) chain with both AAH disorder and nonreciprocal hopping—has not been systematically investigated with topological invariants and entanglement diagnostics. Building on our prior computational survey of generalized SSH models [25], we now investigate this combined system in depth.
@@ III.3 Topological invariants
The biorthogonal polarization [5] $P_{\mathrm{bior}}=\frac{1}{2\pi}\mathrm{Im}[\ln\det F]$, where $F_{mn}=\langle\psi^{L}_{m}|\mathrm{e}^{\mathrm{i}2\pi\hat{X}/N}|\psi^{R}_{n}\rangle$ restricted to occupied states, serves as the real-space topological invariant. The spectral winding number [7] $W(E_{b})=\frac{1}{2\pi\mathrm{i}}\oint_{0}^{2\pi}\mathrm{d}\theta\,\frac{\mathrm{d}}{\mathrm{d}\theta}\ln\det[\hat{H}(\theta)-E_{b}\mathbb{1}]$, with twisted boundary hoppings $w\to w\mathrm{e}^{\pm\mathrm{i}\theta}$, captures the point-gap topology.
@@ V Discussion
Relation to the non-Hermitian AAH model.—The non-Hermitian AAH model (without SSH dimerization) exhibits a single localization–skin-effect phase transition [11, 12, 13]. The SSH sublattice structure qualitatively enriches this picture by introducing: (a) a topological gap that protects the extended phase at small $\lambda$ and $\delta$; (b) two delocalization transitions rather than one, arising from the two-band structure; and (c) a reentrant regime absent in the single-band model, as confirmed by the comparison in Fig. 9. Our results complement those of Ref. [19], which found two delocalization transitions in a non-Hermitian SSH model with quasiperiodic hopping modulation, demonstrating that the SSH sublattice generically produces richer phase structures.

### 31 Mar 2026 v2

@@ I Introduction
The Su–Schrieffer–Heeger (SSH) model [1] is the prototypical one-dimensional topological insulator, supporting symmetry-protected zero-energy edge states when the intercell hopping exceeds the intracell hopping. Adding an Aubry–André–Harper (AAH) on-site potential introduces a metal–insulator transition at a critical modulation strength $\lambda_{c}=2w$, above which all single-particle states become localized [2]. Separately, introducing nonreciprocal hopping renders the Hamiltonian non-Hermitian and gives rise to the non-Hermitian skin effect (NHSE)—the anomalous localization of an extensive number of eigenstates at a single boundary under open boundary conditions [4, 5, 6, 7]. The NHSE is intimately tied to point-gap topology in the complex energy spectrum [8, 9]: the correspondence between skin modes and spectral winding numbers was established in Ref. [7], which further clarified the distinction between trivial and topological skin effects, and the connection to point-gap invariants was developed in parallel in Refs. [6, 8]. The skin effect has also been shown to manifest in non-standard geometries such as fractal lattices, where inner boundaries host distinct skin modes [29].
Each pairwise combination has been studied: the Hermitian SSH–AAH model [18, 19], the non-Hermitian AAH model [12, 13, 14], and the non-Hermitian SSH model without disorder [4, 5]. The non-Hermitian AAH model, in particular, exhibits a competition between skin-effect localization and quasiperiodic localization, with a localization–skin-effect phase transition studied both theoretically [12, 13, 15] and experimentally via photonic quantum walks [17]. Recent work has revealed that quasiperiodicity-induced localization in non-Hermitian systems can proceed sequentially rather than through a single critical potential, giving rise to plateau-like structures with self-similarity [27, 28]. However, the combined model—a dimerized (SSH) chain with both AAH disorder and nonreciprocal hopping—has not been systematically investigated with topological invariants and entanglement diagnostics. Building on our prior computational survey of generalized SSH models [26], we now investigate this combined system in depth.
@@ III.3 Topological invariants
The biorthogonal polarization [5] $P_{\mathrm{bior}}=\frac{1}{2\pi}\mathrm{Im}[\ln\det F]$, where $F_{mn}=\langle\psi^{L}_{m}|\mathrm{e}^{\mathrm{i}2\pi\hat{X}/N}|\psi^{R}_{n}\rangle$ restricted to occupied states, serves as the real-space topological invariant. The spectral winding number [8, 7] $W(E_{b})=\frac{1}{2\pi\mathrm{i}}\oint_{0}^{2\pi}\mathrm{d}\theta\,\frac{\mathrm{d}}{\mathrm{d}\theta}\ln\det[\hat{H}(\theta)-E_{b}\mathbb{1}]$, with twisted boundary hoppings $w\to w\mathrm{e}^{\pm\mathrm{i}\theta}$, captures the point-gap topology.
@@ V Discussion
Relation to the non-Hermitian AAH model.—The non-Hermitian AAH model (without SSH dimerization) exhibits a single localization–skin-effect phase transition [12, 13, 14]. Recent work has shown that this transition can proceed sequentially with self-similar plateau structures rather than through a single critical potential [27, 28]. The SSH sublattice structure qualitatively enriches this picture further by introducing: (a) a topological gap that protects the extended phase at small $\lambda$ and $\delta$; (b) two delocalization transitions rather than one, arising from the two-band structure; and (c) a reentrant regime absent in the single-band model, as confirmed by the comparison in Fig. 9. Our results complement those of Ref. [20], which found two delocalization transitions in a non-Hermitian SSH model with quasiperiodic hopping modulation, demonstrating that the SSH sublattice generically produces richer phase structures.